OpenStudy (anonymous):

Math.. :? Solve. | x -1/5 | =3

OpenStudy (anonymous):

Help me through it... no answers please :)

OpenStudy (anonymous):

x-1/5=3 or x-1/5=-3 x=16/5 x=-14/5

OpenStudy (anonymous):

Thanks :) you didnt really help me through it but thanks anyways haha

OpenStudy (anonymous):

\[\left| a \right|=b\] \[a=b or a=-b\]